### v1

# An Infinite Transitivity Theorem

Abstract.
In this note, we promote an infinite Kadison transitivity theorem on massive $C^{*}$-algebras, including the Calkin algebra. This transitivity stems from the analog of countable degree-1 saturation on pure states which is inherited from these algebras via excision. We show this saturation to be equivalent to several order-theoretic properties on the quantum filter associated to the state, in particular the property of being a quantum P-point. While we show their existence is independent from ZFC, under basic set theoretic assumptions, we produce a plethora of these states. Finally, we find an irreducible representation of the Calkin algebra which fails infinite transitivity.
1. Introduction
1.1. Motivation
The Kadison transitivity theorem is among the most fundamental and classical theorems in $C^{*}$-algebra theory, and is the basis for many results in the representation theory of $C^{*}$-algebras.
Since the advent of continuous model theory in 2008 [YA08], its applications in operator algebras have proliferated. In 2017 [FA17], it was shown by Farah that, in the theory of irreducible representations of $C^{*}$-algebras, ultraproducts admit an infinite version of the Kadison transitivity theorem. Here we show this infinite transitivity also applies to certain irreducible representations on corona algebras $\mathscr{Q}(A)=\mathscr{M}(A)/A$ associated to nonunital $\sigma$-unital $C^{*}$-algebras $A$, in particular, the Calkin algebra $\mathscr{Q}(H)$. These algebras were first shown to have a (not so) weak saturation called countable degree-1 saturation by Farah and Hart in 2011 [FA11].
Around the same time, in 2014 [FA14], Farah and Weaver formalized the notion of maximal quantum filters, the quantum analog of ultrafilters. These filters assisted in the study of excision, which allows one to extract the value of a pure state $|\varphi(b)|$ on an operator $b$ using the quantities $\|aba\|$ for $a$ in the quantum filter associated to $\varphi$. Our version of infinite transitivity relies on a combination of a strong form of excision and countable degree-1 saturation.
1.2. Outline
We say a representation $\pi:C\rightarrow\mathscr{B}(H)$ of a $C^{*}$-algebra $C$ admits infinite transitivity if for every separable subspace $K$ of $H$, the map $C\rightarrow\mathscr{B}(K)$ by $c\mapsto P_{K}\pi(c)P_{K}$ is onto, where $P_{K}$ is the projection onto $K$. This is equivalent to the statement that there is always an operator in $C$ sending a given orthonormal sequence to another given bounded sequence in $H$. The first question one might have is: Does every irreducible representation of the Calkin algebra have infinite transitivity? We produce a negative answer.
Theorem 1.
There exists an irreducible representation of $\mathscr{Q}(H)$ which does not admit infinite transitivity.
Of course, the next question is: Does any representation admit infinite transitivity? Under basic set-theoretic assumptions, in particular the continuum hypothesis, we produce a positive answer.
Theorem 2.
Under the continuum hypothesis, there exists an irreducible representation of $\mathscr{Q}(H)$ which admits infinite transitivity.
To accomplish this, we define the countable degree-$1$ saturation (def. 1.1 and 2.1) of a $C^{*}$-algebra with a state, $(C,\varphi)$, and deduce various consequences of it. In particular, for such a structure with pure $\varphi$, the quantum filter (def. 2.2) associated is a P-filter (def. 2.4). Thanks to the strong excision granted by these quantum filters, we show that many of these order-theoretic consequences of this saturation are, in fact, equivalent to it (theorem 2.8). As such, we call these saturated states P-states. Then we show that the GNS-representation associated to a P-state, called a P-representation, has infinite Kadison transitivity.
Theorem 3.
Let $A$ be a nonunital, $\sigma$-unital $C^{*}$-algebra, and $\pi:\mathscr{Q}(A)\rightarrow\mathscr{B}(H)$ a P-representation. Then $\pi$ admits infinite transitivity. In particular, for every separable subspace $K$ of $H$ and $T\in\mathscr{B}(K)$ self-adjoint (positive, unitary, resp.), there exists $c\in\mathscr{Q}(A)$ such that $\|c\|=\|T\|$, $P_{K}\pi(c)P_{K}=T$, and $c$ is self-adjoint (positive, unitary, resp.).
This follows from infinite Glimm-Kadison on direct sums of P-representations.
Theorem 4.
Let $A$ be a nonunital, $\sigma$-unital $C^{*}$-algebra, and $\pi:\mathscr{Q}(A)\rightarrow\mathscr{B}(H)$ a direct sum of countably many inequivalent P-representations $\pi_{j}:\mathscr{Q}(A)\rightarrow\mathscr{B}(H_{j})$ and $K_{j}$ a separable subspace of $H_{j}$. Then the map $\mathscr{Q}(A)\rightarrow\bigoplus_{j}\mathscr{B}(K_{j})$ by $c\mapsto P_{K}\pi(c)P_{K}$ is onto, where $K=\bigoplus_{j}K_{j}$.
As in theorem 3, we can choose $c$ as we like. With the theory complete, under the existence of a P-point, we produce a P-state, on the corona algebra $\mathscr{Q}(A)$ of a nonunital $A$ with a $\sigma$-unit of projections. However, as implicitly shown in [BI13], there is a forcing in which these states cannot exist.
Finally, we find an irreducible representation of the Calkin algebra which fails infinite transitivity, granting theorem 1. Compare this situation with those of [FA10], wherein the question of whether the commutant of $\mathscr{B}(H)$ in $\mathscr{B}(H)^{\mathscr{U}}$ is trivial also depends on the ultrafilter $\mathscr{U}$. In our case, for a diagonal pure state on $\mathscr{Q}(H)$ (or equivalently $\mathscr{B}(H)$), whether the GNS-representation admits infinite transitivity depends on the corresponding ultrafilter $\mathscr{U}$. Another recent instance of this ultrafilter-dependence phenomenon was in [ZA25].
1.3. Acknowledgements
I would like to thank Tristan Bice and Beatriz Zamora-Avilés for their useful insight into various proofs modified throughout, and Alan Dow for the key to finding a counterexample to infinite transitivity. I give my gratitude to Charles Akemann and David Gao, not only for their help with this paper, but for their overall teaching and mentorship. I also thank Juris Steprāns and Michael Carey for several valuable conversations about set and model theory as I was writing this paper. Of course, I am indebted to Ilijas Farah for his consistent prowess and for teaching me the beauty of combinatorial set theory and massive $C^{*}$-algebras. Without his existing literature, initial prediction, and advisory, this paper would never have come to fruition.
1.4. Preliminaries
Throughout, $A$ denotes an arbitrary $C^{*}$-algebra and $\varphi$ denotes an arbitrary state on $A$. However, when dealing with large $C^{*}$-algebras (like $\mathscr{Q}(A)$) we will denote them by $C$ to avoid confusion with $A$. We denote finite subsets $F\subseteq X$ by $F\Subset X$.
Before understanding saturation of states, we need to understand degree-1 saturation of $C^{*}$-algebras, an extremely powerful property enjoyed by massive corona algebras.
Definition 1.1.
A degree-1 *-polynomial $P(\bar{x})$ over $C$ in variables $\bar{x}=(x_{1},\dots,x_{m})$ is an expression of the form
$P(\bar{x})=\sum_{i\leq m,j\leq n}(a_{ij1}x_{j}a_{ij2}+a_{ij3}x_{j}^{*}a_{ij4})+a_{0},$
where $a_{0}\in C$, $a_{ijk}\in\tilde{C}$, $n\in\mathbb{N}$. A degree-$1$ condition ofCC is an expression of the form
$\|P(\bar{x})\|=r,$
where $P(\bar{x})$ is a degree-$1$ *-polynomial and $r\geq 0$. A countable degree-$1$ type $t(\bar{x})$ is collection of degree-$1$ norm conditions, i.e.
$t(\bar{x})=\{\|P_{j}(\bar{x})\|=r_{j}:j\in\mathbb{N}\}.$
We call $t(\bar{x})$ satisfiable if for every $F\Subset\mathbb{N}$, $\epsilon>0$, there exists $\bar{b}$ in $C$ such that $\max_{j\in F}|\|P_{j}(\bar{b})\|-r_{j}|<\epsilon$. Similarly, $t(\bar{x})$ is realized if there exists $\bar{b}$ such that $\|P_{j}(\bar{b})\|=r_{j}$ for all $j\in\mathbb{N}$. We say that $C$ is countably degree-$1$ saturated if every satisfiable countable degree-$1$ type is realized.
The most exotic examples of countably degree-1 norm saturated $C^{*}$-algebras are those whose saturation does not follow from the standard model-theoretic constructions, namely non-ultraproduct corona algebras. To define coronas, we follow [FA19, 13].
Definition 1.2.
For a nonunital $C^{*}$-algebra $A$, we define the strict topology on $A$ by convergence $a_{\lambda}\rightarrow a$ iff $\lim_{\lambda}\|(a_{\lambda}-a)h\|=\lim_{\lambda}\|h(a_{\lambda}-a)\|=0$ for all $h\in A$. We define the multiplier algebra $\mathscr{M}(A)$ as the completion of $A$ in the strict topology, with algebraic operations defined naturally on the Cauchy classes. Note that this is equivalent to the standard definitions of the multiplier algebra.
The corona algebra $\mathscr{Q}(A)$ is the quotient $C^{*}$-algebra $\mathscr{Q}(A)=\mathscr{M}(A)/A$ with quotient map $\pi:\mathscr{M}(A)\rightarrow\mathscr{Q}(A)$.
When $A$ is nonunital yet $\sigma$-unital, remarkably $\mathscr{Q}(A)$ is countably degree-1 saturated, and a concise proof can be found in [FA19, 15.1.5].
2. Saturated States
Throughout this section, we define countably degree-1 saturated states and work towards understanding them.
Definition 2.1.
[…]
$t(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))=z_{j}:j\in\mathbb{N}\}.$
We define the satisfiability and realizability of $t(\bar{x})$ as in 1.1. We also define the countable degree-1 saturation of $(C,\varphi)$ analogously.
In the remainder of this section, we will work towards the remarkable property of pure states which imbues them with this saturation.
2.1. Quantum Filters
Definition 2.2.
Given a state $\varphi$ on a $C^{*}$-algebra $A$, the quantum filter associated to $\varphi$ is the collection
$\mathscr{F}_{\varphi}=\{a\in A_{+,1}:\varphi(a)=1\},$
where $A_{+,1}$ are the positive, norm 1 elements of $A$. The projection filter associated to $\varphi$ is the collection
$\mathscr{P}_{\varphi}=\{p\in\mathrm{Proj}(A):\varphi(p)=1\}.$
We equip these collections with the order $\ll$ on $A_{+,1}$ by
$a\ll b\text{ iff }ab=a\text{ (equiv. }ba=a).$
Note that, if $p\in\mathrm{Proj}(A)$, $a\in A_{+,1}$, then $p\ll a$ iff $p\leq a$, the usual order on $A_{+,1}$.
There is also an abstract definition of a quantum filter, which is sometimes called a norm filter.
A quantum filter in $A$ is a subset $\mathscr{F}\subseteq A_{+,1}$ with the property that for any $a_{1},\dots,a_{n}\in\mathscr{F}$, $\|a_{1}\cdots a_{n}\|=1$. A maximal quantum filter is a quantum filter not properly contained in any other.
Caution: Quantum filters are not always filters in the order-theoretic sense. More precisely, they are rarely downwards directed under $\ll$ or $\leq$, though interestingly they are always very close [FA19, 5.3.4]. Thankfully, the quantum filters studied in this note will be true (P-)filters under $\ll$.
The beauty of maximal quantum filters is their natural bijective correspondence with pure states, i.e. $\varphi\mapsto\mathscr{F}_{\varphi}$ is a bijection from pure states to maximal quantum filters [FA19, 5.3.8]. While the map $\varphi\mapsto\mathscr{F}_{\varphi}$ is an injection, it is virtually never a bijection, from states to quantum filters. For example, $\{1\}$ is always a quantum filter, but is only of the form $\mathscr{F}_{\varphi}$ when $A=\mathbb{C}$.
2.2. Excision, P-Excision, and Quantum P-points
Quantum filters were first introduced in their relation to excision, a powerful property wherein the value of $|\varphi(b)|$ can be approximated by quantities $\|aba\|$ for $a\in A_{+,1}$.
Definition 2.3.
For a state $\varphi$ on a $C^{*}$-algebra $A$ and subset $\mathscr{F}\subseteq A_{+,1}$, $X\subseteq A$, we say that $\mathscr{F}$ excises $\varphi$ on $X$ if for all $b\in X$,
$\inf_{a\in\mathscr{F}}\|aba-\varphi(b)a^{2}\|=0.$
Most notably, when $\varphi$ is pure, $\mathscr{F}_{\varphi}$ excises $\varphi$ on the entirety of $A$ [FA19, 5.2.1]. However, for small subsets $X\subseteq A$, we can usually excise $\varphi$ using a small portion of $A_{+,1}$. In fact, if $X$ is separable, then there is always some countable $\mathscr{F}\subseteq\mathscr{F}_{\varphi}$ excising $\varphi$ on $X$. This fact was cleverly leveraged in [FA19, 15.3.8] to demonstrate that, when $C$ is countably degree-1 saturated, for every separable $C^{*}$-subalgebra $B$ of $C$, there exists a single $a\in C_{+,1}$ excising $\varphi$ on $B$. This property is quite remarkable, though it has yet to find applications. This is likely due to the fact that $a$ cannot be chosen inside $\mathscr{F}_{\varphi}$, so it is difficult to extract useful information from this single excising element. This is exactly the problem quantum P-points are made to solve, so let us define them now.
Definition 2.4.
A P-filter is a poset $(\mathscr{F},\leq)$ which is upwards closed and $\sigma$-downwards directed. A P-point is an ultrafilter which is a P-filter under $\subseteq^{*}$, inclusion modulo finite sets. A quantum P-point is a maximal quantum filter which is P-filter under $\ll$. That is, for any countable family $\mathscr{F}\subseteq\mathscr{F}_{\varphi}$, there is $a\in\mathscr{F}_{\varphi}$ such that $a\ll b$ for all $b\in\mathscr{F}$. We call $\varphi$ a P-state if $\mathscr{F}_{\varphi}$ is a quantum P-point. Likewise, we call a representation $\pi$ of $C$ a P-representation if it is equivalent to the GNS-representation corresponding to a P-state.
As mentioned above, the most important quality of quantum P-points is their strong excision property. Namely that for every separable $C^{*}$-subalgebra $B$ of $C$, $\varphi$ is excised on $B$ by a single $a\in\mathscr{F}_{\varphi}$. We call this property P-excision. On an intuitive level, it is not hard to see why this would make a serious difference, as $a\in\mathscr{F}_{\varphi}$ allows us to claim $|\varphi(b)|\leq\|aba\|$ for every $b\in C$, not only on the excised subalgebra. This will be crucial when realizing countable degree-1 types in theorem 2.8.
2.3. Equivalences
Let us give some order-theoretic terms relevant to our main equivalence theorem.
Definition 2.5.
A partially ordered set $(\mathscr{F},\leq)$ is countably closed if every decreasing sequence in $\mathscr{F}$ has a lower bound. A subset $\mathscr{P}$ of $\mathscr{F}$ is cofinal if for every $a\in\mathscr{F}$, there exists $b\in\mathscr{P}$ such that $b\leq a$. Lastly, a totally ordered set $(\mathscr{P},<)$ with no endpoints is an $\eta_{1}$-set if for all countable $X,Y\subseteq\mathscr{P}$ such that $a<b$ for all $a\in X$, $b\in Y$, there exists $c\in\mathscr{P}$ s.t. $a<c<b$ for all $a\in X$, $b\in Y$.
In the following lemma to theorem 2.8, we expose a new trick for realizing projections using degree-1 norm types in algebras of real rank zero. Before proceeding, let us define the quantum filter type
$t_{\mathscr{F}}(x)=\{\|x\|=1,x\geq 0,\varphi(x)=1\},$
a finite degree-1 type whose realizations are precisely $\mathscr{F}_{\varphi}$.
Lemma 2.6.
If $C$ has real rank zero, $\varphi$ is a state on $C$, and $(C,\varphi)$ is countably degree-$1$ saturated, then $\mathscr{P}_{\varphi}$ is downwards cofinal in $\mathscr{F}_{\varphi}$.
Proof.
Fix $a\in\mathscr{F}_{\varphi}$ and the type
$t(x)=t_{\mathscr{F}}(x)\cup\{ax=x\}.$
By [FA19, 5.3.10], for each $\epsilon\in(0,1)$, there exists $q\in\mathscr{P}_{\varphi}$ s.t. $q\leq a+\epsilon$, so
$\|aq-q\|^{2}\leq\|(qa-q)^{*}(aq-q)\|=\|q(a-1)^{2}q\|\leq\|qaq-q\|\leq\epsilon.$
Therefore $t(x)$ is satisfiable, so realized by $b\in\mathscr{F}_{\varphi}$ s.t. $b\ll a$. Then we repeat this with $t^{\prime}(x)=t_{\mathscr{F}}(x)\cup\{bx=x\}$ to obtain $c\in\mathscr{F}_{\varphi}$ s.t. $c\ll b\ll a$. Then by real rank zero and [FA19, 2.7.5], there exists a projection $p$ s.t. $c\ll p\ll a$, so $p\in\mathscr{P}_{\varphi}$. ∎
Lemma 2.7.
Let $A$ have real rank zero and $\varphi$ a state on $A$. If $\mathscr{P}_{\varphi}$ is a P-filter, then $\mathscr{F}_{\varphi}$ is a P-filter. In fact, lower bounds of countable families in $\mathscr{F}_{\varphi}$ can be chosen to be projections.
Proof.
Enumerate a sequence $(a_{n})$ in $\mathscr{F}_{\varphi}$. As noted in [FA19, 5.3.10], for all $n,m$, there exists $p_{nm}\in\mathscr{P}_{\varphi}$ s.t. $p_{nm}\leq a_{n}+\frac{1}{m}$. Then by the P-filter property, fix $p\in\mathscr{P}_{\varphi}$, a lower bound for every $p_{nm}$. Then for all $n,m$,
$p\ll p_{nm}\leq a_{n}+\frac{1}{m},$
so $p\ll a_{n}$. Since $(a_{n})$ was arbitrary, we are done. ∎
We state these two lemmas separately from theorem 2.8, as they apply to all states, not only pure ones. Of particular interest to operator algebraists is the fact that these states have weak normality, a property first discovered for projections on the Calkin algebra by Bice in [BI13, 6.2]. That is, decreasing sequences $(a_{n})$ in $A_{+}$ have lower bounds $a\in A_{+}$ with $\varphi(a)=\lim_{n\rightarrow\infty}\varphi(a_{n})$. Also, lemma 2.7 does not require any saturation.
Theorem 2.8.
Let $C$ be countably degree-$1$ saturated and $\varphi$ a pure state on $C$. The following are equivalent:
  1. (1)
    $(C,\varphi)$ is countably degree-1 saturated.
  2. (2)
    $\mathscr{F}_{\varphi}$ is a quantum P-point.
If in addition $C$ has real rank zero, the following are equivalent to $(1),(2)$:
  1. (3)
    $\mathscr{P}_{\varphi}$ is a P-filter.
  2. (4)
    $\mathscr{P}_{\varphi}$ is countably closed.
  3. (5)
    $\mathscr{P}_{\varphi}$ is cofinal in $\mathscr{F}_{\varphi}$.
If $C$ also has no minimal projections, the following is equivalent to (1)(1)-$(5)$:
  1. (6)
    Maximal chains in $\mathscr{P}_{\varphi}\setminus\{1\}$ are $\eta_{1}$-sets.
If $C$ is also purely infinite and simple, the following is equivalent to (1)(1)-$(6)$:
  1. (7)
    $\mathscr{P}_{\varphi}$ is a filter.
1234567
Figure 1. Equivalences
Proof.
Figure 1 depicts the layout of equivalences. Firstly, note that $(3)\Leftrightarrow(4)\wedge(7)$, denoted by dashed arrows in figure 1, regardless of the assumptions on $C$, $(1)\Rightarrow(5)$ is the statement of 2.6, and $(3)\Rightarrow(2)$ is the statement of 2.7. We will prove the other directions in the following order:
$(1)\Rightarrow(2),(3);(4)\Leftrightarrow(7);(5)\Rightarrow(4);(4)\Leftrightarrow(6)$
and we save $(2)\Rightarrow(1)$ for later, denoted by the dotted arrow in figure 1.
For $(1)\Rightarrow(2)$, let $Y\subseteq\mathscr{F}_{\varphi}$ be countable and define the type
$t(x)=t_{\mathscr{F}}(x)\cup\{ax=x:a\in Y\}.$
By [FA19, 5.3.4], the set $\mathscr{D}=\{a\in C_{+,1}:1-a\in\Lambda\}$ is dense in $\mathscr{F}_{\varphi}$, where
$\Lambda=\{b\in(L_{\varphi}\cap L_{\varphi}^{*})_{+}:b<1\}.$
Now define an approximate unit $(e_{\gamma})$ for $L_{\varphi}\cap L_{\varphi}^{*}$. Notice that for each $a\in\mathscr{D}$, $(1-e_{\gamma})a-(1-e_{\gamma})=(e_{\gamma}-1)(1-a)=e_{\gamma}(1-a)-(1-a).$ Now fix $F\Subset Y$, $\epsilon>0$, and for each $a\in F$, choose $a^{\prime}\in\mathscr{D}$ s.t. $\|a^{\prime}-a\|<\frac{\epsilon}{2}$. There exists $\gamma$ s.t. for all $a\in F$, $\|e_{\gamma}(1-a^{\prime})-(1-a^{\prime})\|<\frac{\epsilon}{2}$. Therefore
$\|(1-e_{\gamma})a-(1-e_{\gamma})\|$ $\leq\|1-e_{\gamma}\|\|a-a^{\prime}\|+\|e_{\gamma}a^{\prime}-a^{\prime}\|$
$<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon$
so $(1-e_{\gamma})$ satisfies $t(x)$, so it is realized by $b\in\mathscr{F}_{\varphi}$ s.t. $b\ll a$ for all $a\in Y$. Since $\mathscr{F}$ was arbitrary, we are done.
For $(1)\Rightarrow(3)$, fix countable $Y\subseteq\mathscr{P}_{\varphi}\subseteq\mathscr{F}_{\varphi}$. Then by (2), fix a lower bound $b\in\mathscr{F}_{\varphi}$ for $Y$. Then by lemma 2.6 there exists $p\in\mathscr{P}_{\varphi}$ s.t. $p\ll b$, so $p$ a lower bound for $Y$.
For $(4)\Leftrightarrow(7)$, we apply both directions of [BI13, 4.1], so $(4)\vee(7)\Rightarrow(4)\wedge(7)\Rightarrow(3)$.
For $(5)\Rightarrow(4)$, enumerate $(p_{n})$ in $\mathscr{P}_{\varphi}$. Then define $a\in C_{+,1}$ by
$a=\sum_{n=1}^{\infty}\frac{1}{n(n+1)}p_{n}$
and notice
$\varphi(a)=\sum_{n=1}^{\infty}\frac{1}{n(n+1)}=1,$
so $a\in\mathscr{F}_{\varphi}$. By cofinality, fix some $p\in\mathscr{P}_{\varphi}$ s.t. $p\ll a$. Then
$p\leq a$ $=\sum_{n=1}^{\infty}\frac{1}{n(n+1)}p_{n}$
$\leq\sum_{n=1}^{m}\frac{1}{n(n+1)}+\sum_{n=m+1}^{\infty}\frac{1}{n(n+1)}p_{n}$
$\leq\bigg(1-\frac{1}{m}\bigg)+\frac{1}{m}p_{m}$
$\leq p_{m}+\bigg(1-\frac{1}{m}\bigg)(1-p_{m})$
Since $1-\frac{1}{m}<1$, this is only possible if $p\leq p_{m}$, so $p\ll p_{m}$. Since $m$ and $(p_{n})$ were arbitrary, $\mathscr{P}_{\varphi}$ is countably closed.
[…]
Compare condition (6) with [HA98, 3], which states that maximal chains of projections in $\mathscr{Q}(H)$ are $\eta_{1}$-sets. By the above reasoning, this theorem also holds for any countably degree-1 saturated $C$ with with real rank zero and no minimal projections.
Corollary 2.9.
Let $\varphi$ be a pure state s.t. $(C,\varphi)$ countably degree-$1$ saturated. Then $\varphi$ has P-excision. If in addition, $C$ has real rank zero, then the excising element can always be chosen to be a projection.
Proof.
Let $B$ be a separable $C^{*}$-subalgebra of $C$. As noted in [FA19, 15.3.7], Löwenheim-Skolem grants us a countable $Y\subseteq\mathscr{F}_{\varphi}$ which excises $\varphi$ on $B$. By theorem 2.8 $(1)\Rightarrow(2)$, this subset has a lower bound $a\in\mathscr{F}_{\varphi}$. Then as noted in [FA19, 15.3.8], for $b\in B$, $c\in Y$,
$\|aba-\varphi(b)a^{2}\|=\|a(b-\varphi(b))a\|=\|ac(b-\varphi(b))ca\|\leq\|a\|^{2}\|c(b-\varphi(b))c\|,$
so
$\|aba-\varphi(b)a^{2}\|\leq\|a\|^{2}\inf_{c\in\mathscr{F}_{\varphi}}\|cbc-\varphi(b)c^{2}\|=0.$
Since $b$ was arbitrary, we are done. ∎
This P-excision proof actually gives us the crucial tool for proving $(2)\Rightarrow(1)$ of theorem 2.8.
Proof of Theorem 2.8.
For $(2)\Rightarrow(1)$, fix a satisfiable degree-1 type
$t(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))=z_{j}:j\in\mathbb{N}\}.$
Without loss of generality, by the linearity of $\varphi$, whenever $\rho_{j}=\varphi(\cdot)$, assume $z_{j}=0$.
Now fix, for each $m\in\mathbb{N}$, $\bar{b}_{m}$ s.t. $\max_{j\leq m}|\rho_{j}(P_{j}(\bar{b}_{m}))-z_{j}|<\frac{1}{m}.$ Now let $X$ be the collection of entries of each $\bar{b}_{m}$ and coefficients of each $P_{j}$. Then $X$ is countable, so $B=C^{*}(X)$ is separable. Mimicking the proof of 2.9, there exists a single $a\in\mathscr{F}_{\varphi}$ which excises $\varphi$ on $B$. Now define the modified predicate
$\rho^{\prime}_{j}=\begin{cases}\|\cdot\|,&\rho_{j}=\|\cdot\|;\\ \|a\cdot a\|,&\rho_{j}=\varphi(\cdot);\end{cases}$
fix the degree-$1$ norm type
$t^{\prime}(\bar{x})=\{\rho^{\prime}_{j}(P_{j}(\bar{x}))=z_{j}:j\in\mathbb{N}\}$
and notice $(\bar{b}_{m})$ satisfies $t^{\prime}(\bar{x})$. Since $C$ is countably degree-1 saturated, $t^{\prime}(\bar{x})$ is realized. When $\rho_{j}=\varphi(\cdot)$, we have
$|\varphi(P_{j}(\bar{b}))|=|\varphi(aP_{j}(\bar{b})a)|\leq\|aP_{j}(\bar{b})a\|=0.$
Therefore, $\bar{b}$ realizes $t(\bar{x})$. ∎
Now that we have these equivalences, we will simply refer to pure $\varphi$ such that $(C,\varphi)$ is countably degree-1 saturated as a P-state.
3. Infinite Transitivity
The reader may have noticed a concerning gap between the countable degree-1 saturation of $(C,\varphi)$ and an infinite Kadison transitivity theorem: Vector equations in $H_{\varphi}$ are degree-2 conditions of the form
$0=\|xa-b\|^{\varphi}_{2}=\varphi((xa-b)^{*}(xa-b))^{\frac{1}{2}}.$
The following lemma shows the true potential of P-states and addresses these concerns.
Lemma 3.1.
Let $C$ be countably degree-1 saturated. Then $C$ is saturated over countable types consisting of the following conditions
  1. (1)
    $\|P(\bar{x})\|\in K$
  2. (2)
    $\|P(\bar{x})\|^{\varphi}_{2}\leq r$
  3. (3)
    $\varphi(P(\bar{x}))\in\Lambda$
where $P(\bar{x})$ a degree-$1$ *-polynomial, $K\subseteq[0,\infty)$, $\Lambda\subseteq\mathbb{C}$ both compact, $r\geq 0$, and $\varphi$ a P-state.
Proof.
Let $t(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))\in K_{j}:j\in\mathbb{N}\}$ be a satisfiable countable type consisting of conditions above. To prove this, we will perform two conversions of the above
Since $t(\bar{x})$ is satisfiable, fix $\bar{b}_{m}$ s.t. $\max_{j\leq m}d(\rho_{j}(P_{j}(\bar{b}_{m})),K_{j})\leq\frac{1}{m}$. Then fix an ultrafilter $\mathscr{U}$ on $\mathbb{N}$. Let $z_{j}=\lim_{m\rightarrow\mathscr{U}}\rho_{j}(P_{j}(\bar{b}_{m}))$. Define
$\sim_{j}=\begin{cases}\leq,&\rho_{j}=\|\cdot\|_{2}^{\varphi};\\ =,&\text{otherwise}.\end{cases}$
Then
$t^{\prime}(\bar{x})=\{\rho_{j}(P_{j}(\bar{x}))\sim_{j}z_{j}:j\in\mathbb{N}\}$
is satisfiable type whose realizations realize $t(\bar{x})$. Now for each $j$ s.t. $\rho_{j}=\varphi$ or $\|\cdot\|_{2}^{\varphi}$, choose $a_{\varphi}\in\mathscr{F}_{\varphi}$ excising the $C^{*}$-subalgebra $B$ generated by coefficients of $P_{j}(\bar{x})$ and entries of $\bar{b}_{m}$. Then, for each $j$ s.t. $\rho_{j}=\varphi$, choose $d_{\varphi}\in B$ s.t. $\varphi(d_{\varphi})=z_{j}$. Define
$P^{\prime}_{j}(\bar{x})=\begin{cases}P_{j}(\bar{x})-d_{\varphi},&\rho_{j}=\varphi;\\ P_{j}(\bar{x}),&\text{otherwise}.\end{cases}$
$\rho^{\prime}_{j}=\begin{cases}\|a_{\varphi}\cdot a_{\varphi}\|,&\rho_{j}=\varphi;\\ \|\cdot a_{\varphi}\|,&\rho_{j}=\|\cdot\|^{\varphi}_{2};\\ \rho_{j},&\text{otherwise}.\end{cases}$
$r_{j}=\begin{cases}0,&\rho_{j}=\varphi;\\ z_{j},&\text{otherwise}.\end{cases}$
The type
$t^{\prime\prime}(\bar{x})=\{\rho^{\prime}_{j}(P^{\prime}_{j}(\bar{x}))\sim_{j}r_{j}:j\in\mathbb{N}\}$
is a satisfiable generalized countable degree-1 norm type whose realizations realize $t^{\prime}(\bar{x})$, thus realize $t(\bar{x})$. Therefore, by [FA19, 15.2.10], $t(\bar{x})$ is realized. ∎
We now prove the main results of the paper, infinite Glimm-Kadison for P-representations on coronas of countably degree-1 saturated $C^{*}$-algebras, from which infinite transitivity will follow.
Proof of Theorem 4.
Let $K_{j}$ be a separable subspace of $H_{j}$ and numerate an orthonormal basis $(\xi_{j})$ of $K=\bigoplus_{j}K_{j}$ compatible with each $K_{j}$. Every $T\in\bigoplus_{j}\mathscr{B}(K_{j})$ is determined by $(T\xi_{j})$. For each $j$, there exists some P-state $\varphi_{j}$ such that there is a lift $a_{j}+L_{\varphi_{j}}=\xi_{j}$. Similarly, there is some lift $b_{j}+L_{\varphi_{j}}=T\xi_{j}$. Define the type
$t(x)=\{\|x\|=\|T\|,\|xa_{j}-b_{j}\|^{\varphi_{j}}_{2}=0:j\in\mathbb{N}\}.$
Since $t(x)$ consists of conditions from 3.1, it suffices to show that it is satisfiable. Fix $F\Subset\mathbb{N}.$ Because $\{\xi_{j}:j\in F\}$ is orthogonal, by the Glimm-Kadison theorem, there exists $d\in C$ s.t. $\pi(d)\xi_{j}=\eta_{j}$ for all $j\in F.$ That is, $\|da_{j}-b_{j}\|^{\varphi}_{2}=0,$ for all $j\in F.$ Note that, because the vectors are uniformly bounded, all these partial realizations can be taken to be in some fixed sort. Therefore $t(x)$ is satisfiable and is realized by $c\in C$, the promised element.
For self-adjoint or positive, add the necessary generalized degree-$1$ norm conditions to $t(x)$ [FA19, 15.2.9], yielding a satisfiable type which is realized by countably degree-1 saturated $C$, thanks to 3.1.
For unitary, notice that by the Borel functional calculus, $T$ has self-adjoint complex logarithm $S\in\bigoplus_{j}\mathscr{B}(K_{j})$. Then, by the self-adjoint case, there exists self-adjoint $s\in C$ so that $P_{K}\pi(s)P_{K}=S$, so $P_{K}\pi(e^{is})P_{K}=T$, so $c=e^{is}$ is our desired unitary. ∎
We note that we can readily produce infinite versions of [FA19, 3.5.5, 3.8.3, 3.9.6, 5.1.3, 5.6.6] in the case of P-states/representations, but from here these proofs are so elementary that they would only serve to dilute this text.
4. Existence of P-states
The conclusion of theorem 2.8 can be leveraged to show that, in the right model of ZFC, there are no P-states on many corona algebras. However, in any model of ZFC which admits a P-point (an extremely weak assumption), we can produce a P-state on these coronas.
Definition 4.1.
Given nonunital $A$ with a $\sigma$-unit of projections $(e_{n})$, define the map $p:\mathscr{P}(\mathbb{N})\rightarrow\mathrm{Proj}(\mathscr{M}(A))$ by $p_{X}$ equal to the strict limit of the sequence
$p_{X,m}=\sum_{n\in X,n\leq m}f_{n}$
where $f_{n}=e_{n}-e_{n-1}$, $e_{-1}=0$. For each $\varphi\in S(\mathscr{Q}(A))$, define
$\mathscr{X}_{\varphi}=\{X\subseteq\mathbb{N}:\varphi(\pi(p_{X}))=1\}.$
The following lemma was originally proven for the Calkin algebra by Zamora-Aviles [ZA10, 2.5.10]. Here we provide a version generalized to coronas with $\sigma$-units of projections.
Lemma 4.2.
Let $A$ be a nonunital $\sigma$-unital $C^{*}$-algebra and $(e_{n})$ a $\sigma$-unit of projections. For any decreasing sequence $X_{n}\subseteq\mathbb{N}$, $a\leq\pi(p_{X_{n}})$ for all $n$ if and only if there exists a pseudointersection $X\subseteq\mathbb{N}$ of $(X_{n})$ s.t. $a\leq\pi(p_{X})$.
Proof.
Without loss of generality, suppose $X_{1}=\mathbb{N}$ and $X_{n}\setminus X_{n+1}$ infinite. Enumerate $X_{n}\setminus X_{n+1}=\{j^{n}_{i}:i\in\mathbb{N}\}$. Now define $p^{n}:\mathscr{P}(\mathbb{N})\rightarrow\mathrm{Proj}(\mathscr{M}(A))$ by $p^{n}_{Y}=p_{\{j^{n}_{i}:i\in Y\}}$. Since $a=\pi(\tilde{a})\leq\pi(p_{X_{n}})$, we have
$\pi(p_{X_{n}\setminus X_{n+1}}\tilde{a})=(\pi(p_{X_{n}})-\pi(p_{X_{n+1}}))a=a-a=0,$
so $\lim_{k\rightarrow\infty}\|(1-e_{k})p_{X_{n}\setminus X_{n+1}}\tilde{a}\|=\lim_{k\rightarrow\infty}\|p^{n}_{\mathbb{N}\setminus k}\tilde{a}\|=0,$ so the function $f_{\epsilon}\in\mathbb{N}^{\mathbb{N}}$ by
$f_{\epsilon}(n)=\min\Big\{k:\|p^{n}_{\mathbb{N}\setminus k}\tilde{a}\|<\frac{\epsilon}{2^{n+1}}\Big\}$
is well-defined. Fix $f\in\mathbb{N}^{\mathbb{N}}$ s.t. $f_{\frac{1}{m}}\leq^{*}f$ for all $m$. That is, $f_{\frac{1}{m}}(n)\leq f(n)$ for e.a. $n$. Now we claim that $X=\bigcup_{n\in\mathbb{N}}\{j^{n}_{i}:i\leq f(n)\}$ is our desired pseudointersection. Since $X_{n}$ decreases from $\mathbb{N}$, $X_{n}\setminus X_{n+1}$ partitions $\mathbb{N}$, so
$X\setminus X_{n}=\bigcup_{j=1}^{n-1}X\cap(X_{j}\setminus X_{j+1})=\bigcup_{j=1}^{n-1}\{j^{n}_{i}:i\leq f(n)\},$
so
$|X\setminus X_{n}|=\sum_{j=1}^{n-1}(f(j)-1)<\infty,$
thus $X$ is a pseudointersection. Now we claim $a\leq\pi(p_{X})$, and it suffices to show that $\|(1-e_{k})p_{\mathbb{N}\setminus X}\tilde{a}\|\rightarrow 0$. By $f_{\frac{1}{m}}\leq^{*}f$, we have
$\|(1-e_{k})p_{\mathbb{N}\setminus X}\tilde{a}\|$ $\leq\sum_{n=1}^{\infty}\|(1-e_{k})p^{n}_{\mathbb{N}\setminus f(n)}\tilde{a}\|$
$\leq\sum_{n=1}^{N-1}\|(1-e_{k})p^{n}_{\mathbb{N}\setminus f(n)}\tilde{a}\|+\sum_{n=N}^{\infty}\frac{1}{m2^{n+1}}$
$\leq\sum_{n=1}^{N-1}\|(1-e_{k})p^{n}_{\mathbb{N}\setminus f(n)}\tilde{a}\|+\frac{1}{m2^{N}}$
$\leq\sum_{n=1}^{N-1}\|(1-e_{k})p_{X_{n}\setminus X_{n+1}}\tilde{a}\|+\frac{1}{m2^{N}}$
$\leq\frac{1}{m}+\frac{1}{m}\leq\frac{2}{m}$
by taking $k$ so that $\|(1-e_{k})p_{X_{n}\setminus X_{n+1}}\tilde{a}\|\leq\frac{1}{(N-1)m}$ for all $n<N$. ∎
The following theorem gives us a method for finding P-states.
Theorem 4.3.
Let $A$ be a nonunital $\sigma$-unital $C^{*}$-algebra. For every $\sigma$-unit of projections $(e_{n})$ of $A$, there exists $\varphi:\beta\mathbb{N}\setminus\mathbb{N}\hookrightarrow P(\mathscr{Q}(A))$ such that $\mathscr{X}_{\varphi}=\mathscr{U}$ and $\varphi_{\mathscr{U}}$ is a P-state if and only if $\mathscr{U}$ is a P-point.
Proof.
We begin by fixing $\varphi_{n}\in P(A)$ s.t. $\varphi_{n}(f_{n})=1$. Since $(e_{n})$ strictly increasing and consists of projections, $\|f_{n}\|=1$ so such a state always exists. By [MU90, 3.3.9], these each have unique extensions $\tilde{\varphi_{n}}\in P(\mathscr{M}(A))$. Now define $\varphi:\beta\mathbb{N}\setminus\mathbb{N}\hookrightarrow S(\mathscr{Q}(A))$ by
$\varphi_{\mathscr{U}}(\pi(a))=\lim_{n\rightarrow\mathscr{U}}\tilde{\varphi_{n}}(a).$
We know these are states because they lift to weak-limits of pure states on $\mathscr{M}(A)$ and they are pure by [AN79, 1].
By theorem 2.8, it suffices to show that $\mathscr{F}_{\varphi_{\mathscr{U}}}$ is a quantum P-point iff $\mathscr{U}$ is a P-point. Suppose that $\mathscr{U}$ is a P-point and fix a sequence $(a_{n})$ in $\mathscr{F}_{\varphi}$ with respective lifts $\tilde{a_{n}}$. For each $m,n$, define
$X_{m,n}=\Big\{k\in\mathbb{N}:\tilde{\varphi_{k}}(\tilde{a_{n}})>1-\frac{1}{m}\Big\}.$
By construction, $X_{m,n}\in\mathscr{U}$. Since $\mathscr{U}$ is a P-point, there exists $X\in\mathscr{U}$ s.t. for all $m,n$, $X\setminus X_{m,n}$ finite. Enumerate $X=\{k_{j}:j\in\mathbb{N}\}$ so that for all $m,n$, $\tilde{\varphi_{k_{j}}}(\tilde{a_{n}})\geq 1-\frac{1}{m}$ for eventually all $j$. Therefore $\lim_{j\rightarrow\infty}\tilde{\varphi_{k_{j}}}(\tilde{a_{n}})=1$. Our claim is that $a=\pi(p_{X})$ is a lower bound for $(a_{n})$. Since $X\in\mathscr{U}$, $\varphi_{\mathscr{U}}(\pi(p_{X}))=1$, so $a\in\mathscr{F}_{\varphi}$. Now notice that
$\|\pi((\tilde{a_{n}}-1)p_{Y})\|=\lim_{k\rightarrow\infty}\|(\tilde{a_{n}}-1)p_{X}(1-e_{k})\|=\|(\tilde{a_{n}}-1)(p_{X}-p_{X})\|=0,$
since $p_{X}e_{n}\rightarrow p_{X}$ strictly. Therefore
$a_{n}a-a=\pi(\tilde{a_{n}}p_{X}-p_{X})=\pi((\tilde{a_{n}}-1)p_{X})=0,$
i.e. $a\ll a_{n}$ for all $n$.
Now suppose $\mathscr{U}$ is not a P-point, and fix decreasing $X_{n}\in\mathscr{U}$ with no pseudointersection in $\mathscr{U}$. Our claim is that $\pi(p_{X_{n}})$ has no lower bound in $\mathscr{P}_{\varphi_{\mathscr{U}}}$. By 4.2, any lower bound $a\in\mathscr{F}_{\varphi}$ of $(\pi(p_{X_{n}}))$ must in fact have $a\ll\pi(p_{X})$, where $X$ is a pseudointersection of $(X_{n})$, therefore no such $a$ exists. ∎
The existence of a P-point is extremely common in models of ZFC. In fact, until recently [CH19], there were very few known models of ZFC which have no P-points. Their existence is implied by the continuum hypothesis, is consistent with various standard forcing axioms, and even $\text{OCA}_{\text{T}}$. For the operator algebraist reader who does not fancy the continuum hypothesis, rest assured that P-states are likely relatively consistent which whichever axioms one fancies most.
Lemma 4.4.
Let $A$ be a nonunital $C^{*}$-algebra with a $\sigma$-unit of projections, and $\varphi\in P(\mathscr{Q}(A))$. If $\varphi$ is a P-state, then $\mathscr{X}_{\varphi}$ is a non-meagre P-filter. Also, for any decreasing $X_{n}\subseteq\mathbb{N}$, there exists a pseudointersection $X$ of $(X_{n})$ s.t. $\varphi(\pi(p_{X}))=\inf_{n}\varphi(\pi(p_{X_{n}}))$. Also, if $\varphi(\pi(p_{X_{n}}))\not\rightarrow 0$, there exists $(k_{n})$ in $\mathbb{N}$ increasing and $X\subseteq\mathbb{N}$ s.t. $X\subseteq^{*}\bigcup_{j\in[k_{n},k_{n+1})}X_{j}$ and $\varphi(X)>0$.
Proof.
We mimic the proof of [BI13, 6.2]. To show that $\mathscr{X}_{\varphi}$ is non-meagre, take an arbitrary interval partition $(I_{n})$ of $\mathbb{N}$ and let $B$ be the $A$-strict completion of the $C^{*}$-subalgebra generated by $\{p_{I_{n}}:n\in\mathbb{N}\}$. This is isomorphic to $\ell^{\infty}$ and its quotient $B$ is isomorphic to $\ell^{\infty}/c_{0}$. Then $\mathrm{Proj}(B)$ is countably closed and atomless, so by [BI13, 6.1], $\mathscr{P}_{\varphi}^{B}\neq\{1\}$. Therefore, $\mathscr{X}_{\varphi}$ is non-meagre.
Using the obvious type, there exists $a\in\mathscr{P}_{\varphi}$ s.t. $a\ll\pi(p_{X_{n}})$ and $\varphi(a)=\inf_{n}\varphi(\pi(p_{X_{n}}))$. Then by 4.2, there is a pseudointersection $X$ of $(X_{n})$ s.t. $a\ll\pi(p_{X})$, so $1=\varphi(a)\leq\varphi(\pi(p_{X}))$. It then follows that $\mathscr{X}_{\varphi}$ is a P-filter.
For the last part of the claim, let $\epsilon=\limsup_{n\rightarrow\infty}\frac{\varphi(\pi(p_{X_{n}}))}{2}>0$. If
$\varphi(\pi(p_{X_{m_{i+1}}\setminus\bigcup_{j\in[k,m_{i})}X_{j}}))\geq\frac{\epsilon}{3^{n+1}}$
for $i<l$, where $k<m_{0}<\cdots<m_{l}$, then $l\epsilon<3^{n+1}$. We may then choose $k_{n}$ s.t. $\varphi(\pi(p_{X_{j}}))>\epsilon$ for some $k_{n}\leq j<k_{n+1}$ and
$\varphi(\pi(p_{X_{m}\setminus\bigcup_{j\in[k_{n},k_{n+1})}X_{j}}))<\frac{\epsilon}{3^{n+1}}$
for all $m\geq k_{n+1}$. Then
$\varphi(\pi(p_{\bigcap_{n\leq m}\bigcup_{j\in[k_{n},k_{n+1})}X_{j}}))>\epsilon-\sum_{n<m}\frac{\epsilon}{3^{n+1}}\geq\epsilon-\frac{\epsilon}{2}=\frac{\epsilon}{2}$
Then by normality, we have $X\subseteq^{*}\bigcup_{j\in[k_{n},k_{n+1})}X_{j}$ so that $\varphi(\pi(p_{X}))\geq\frac{\epsilon}{2}>0$. ∎
We now have enough to produce P-states.
Theorem 4.5.
Both of the following statements are relatively consistent with ZFC.
  1. (1)
    For every nonunital $A$ with a $\sigma$-unit of projections, there exist $2^{\mathfrak{c}}$-many inequivalent P-states $\varphi$ on $\mathscr{Q}(A)$.
  2. (2)
    For every nonunital $A$ with a $\sigma$-unit of projections, there exist no P-states on $\mathscr{Q}(A)$.
Proof.
For (1), under the continuum hypothesis, there are $2^{\mathfrak{c}}$ RK-inequivalent selective ultrafilters (which are P-points) on $\mathbb{N}$, so the map $\varphi:\beta\mathbb{N}\setminus\mathbb{N}\hookrightarrow P(\mathscr{Q}(A))$ in 4.3 produces $2^{\mathfrak{c}}$ inequivalent P-states.
For (2), we use the forcing argument outlined in lemma [BI13, 6.5], where we use 4.4 in place of [BI13, 6.2]. In this model, there are no $\varphi$ in $\mathscr{Q}(A)$ having the conclusion of 4.4, so there are no P-states. ∎
4.1. Counterexample to Infinite Transitivity
The question of whether infinite transitivity in $(\mathscr{Q}(H),\pi_{\varphi})$ implies degree-1 saturation of $(\mathscr{Q}(H),\varphi)$ is fairly opaque. In this section, we at least prove that in ZFC, many massive coronas have an irreducible representation which does not admit infinite transitivity.
Definition 4.6.
For $A\subseteq\mathbb{N}$, $j\in\mathbb{N}$, we define $X+j=\{m+j:m\in X\}$ and for an ultrafilter $\mathscr{U}$ on $\mathbb{N}$, $\mathscr{U}+j=\{X+j:X\in\mathscr{U}\}$. Since $\beta\mathbb{N}$ is compact, we can define $\mathscr{U}+\mathscr{V}=\lim_{j\rightarrow\mathscr{V}}(\mathscr{U}+j)$. We call a nonprincipal ultrafilter $\mathscr{U}$ idempotent if $\mathscr{U}+\mathscr{U}=\mathscr{U}$.
By [EL58, 1], every compact topological space with a semigroup structure whose left operation is continuous admits an idempotent element. The addition defined in 4.6 plainly grants this semigroup structure on $\beta\mathbb{N}$, so there is $\mathscr{U}\in\beta\mathbb{N}$ idempotent, and this cannot be principal.
Proof of Theorem 1.
Since we already proved the existence of $P$-states on $\mathscr{Q}(H)$ under the continuum hypothesis, it suffices to show there is an irreducible representation of $\mathscr{Q}(H)$ which fails infinite transitivity. In fact, such a representation exists in ZFC.
Let $\mathscr{U}$ be an idempotent ultrafilter on $\mathbb{N}$. Let $s\in\mathscr{B}(H)$ be the right shift along an orthonormal basis $(\eta_{n})$ of $H$ and $\varphi_{\mathscr{U}}$ as in 4.3 along $(\eta_{n})$. Let $\xi_{j}=\pi(s^{j})+L_{\varphi_{\mathscr{U}}}$. Then the equations
$\pi_{\varphi}(a)\xi_{j}=\delta_{j0}\xi_{j}$
imply
$\varphi(\pi(s^{j})^{*}a\pi(s^{j}))=\lim_{n\rightarrow\mathscr{U}}a_{n-j}=\delta_{j0}$
where $a_{n}=\langle a\eta_{n},\eta_{n}\rangle$. Thus $X_{1}=\{n\in\mathbb{N}:|a_{n}-1|<\epsilon\}\in\mathscr{U}$ and for all $j>1$, $X_{j}=\{n\geq j:|a_{n-j}|<\epsilon\}\in\mathscr{U}$. However, $X_{j}=(X_{1}+j)\setminus j$, so by idempotency, $\{j\in\mathbb{N}:X_{1}+j\in\mathscr{U}\}\in\mathscr{U}$, so $\varphi(\pi(s^{j})^{*}a\pi(s^{j}))=1$ for some $j\in\mathbb{N}$. Therefore, there is no $a\in\mathscr{Q}(H)$ s.t. $\pi_{\varphi}(a)\xi_{j}=\delta_{j0}\xi_{j}$, so infinite transitivity fails. ∎
This tells us that infinite transitivity is not a property intrinsic to all irreducible representations of countably degree-1 saturated $C^{*}$-algebras, so the infinite transitivity of P-representations comes from the nontrivial interplay between the state and the algebra.